# Toward a test of Gaussianity of a gravitational wave background

### v1

@@ I Introduction
The first direct detection of a gravitational wave (GW) has ushered in a new era in astronomy Abbott et al. (2016). Hundreds of GW events have now been observed from Solar mass compact objects Abbott et al. (2023), each one carrying important information about the strong gravity regime Abbott et al. (2021). The next breakthrough in GW astronomy is the detection of a gravitational wave background (GWB) Caprini and Figueroa (2018); Christensen (2019); Romano (2019); Moore and Vecchio (2021); Pol et al. (2021, 2022); Staelens and Nelemans (2024), coming from a superposition of GWs from sources that are too weak to be resolved individually. Efforts are ongoing in order to resolve the GWB in the sub-kiloherz band using ground-based GW detectors Abbott et al. (2009); Shoemaker (2019); Lehoucq et al. (2023) and in the millihertz band in future space-based GW observations Bartolo et al. (2022); Cheng et al. (2022); Liang et al. (2022); Muratore et al. (2024). Recently, pulsar timing array (PTA) collaborations have reported a compelling evidence in favor of the presence of a nanohertz stochastic GWB Agazie et al. (2023a); Reardon et al. (2023); Antoniadis et al. (2023); Xu et al. (2023), consistent with a common-spectrum process across pulsars with the signature Hellings and Downs (HD) correlation Hellings and Downs (1983). This signal is traditionally interpreted as coming from a population of supermassive black hole binaries (SMBHB) Sazhin (1978); Detweiler (1979); Phinney (2001); Wyithe and Loeb (2003); Sesana et al. (2004, 2008); Burke-Spolaor et al. (2019); Sato-Polito et al. (2023); Sato-Polito and Kamionkowski (2024); Bi et al. (2023); Sato-Polito and Zaldarriaga (2024); Raidal et al. (2024). However, the PTAs’ observation also allows for an interesting set of sources of the nanohertz GWB, rooted cosmologically Chen et al. (2020); Ellis and Lewicki (2021); Arzoumanian et al. (2021a); Buchmuller et al. (2021); Xue et al. (2021); Sharma (2022); Figueroa et al. (2024); Ellis et al. (2024a); Saeedzadeh et al. (2024); Winkler and Freese (2024). A challenge to the field now is to identify the signatures in the observation that can distinguish the different sources of a GWB, ultimately settling the issue of whether an observed GWB signal is astrophysical or cosmological.
@@
Different sources of a GWB produce the same HD spatial correlation, only with a different spectrum. A logical step to determine the source of the the signal is thus by calculating the expected spectrum for each source and computing the respective evidences relative to the data. This approach has been widely adopted by the PTA community, but has not ruled out any significant possibilities, given the present data Afzal et al. (2023); Ellis et al. (2024b). The field is optimistic that forthcoming iterations of the data, with more and longer timed pulsars, may eventually enable PTAs to contribute to the source debate. This poses the question whether there are further untapped aspects of the observation that may be able to help in out distinguishing the source of an observed GWB signal, not only in PTA but in general. This brings us to the theme of this work which is testing the Gaussianity of a GWB.
@@
In this work, we bare down signatures of the GWB signal in the one- and two-point functions’ cumulants, referring to PTA simulations Babak et al. (2024) as a proof-of-concept. In particular, a Gaussian GWB requires that all the higher-point information and the cumulants of the observable reduce to products of the power spectrum, and the spatial correlation. The relation this has on the source is that the level of Gaussianity that is expected to be carried by the signal is different for astrophysical/SMBHBs and cosmological sources, with the latter expected to display a higher degree of Gaussianity. Testing the Gaussianity of the signal thus boils down to measuring the cumulants of a relevant observable, and checking if there is additional information in the cumulants that is not already present in the power spectrum.
@@
This direction we embark on is reminiscent of tests of isotropy and statistics in the cosmic microwave background (CMB) Akrami et al. (2020). In CMB analysis, the one-point information is constrained to assess if the data shows any significant departures from Gaussianity. If there is, then a primordial theory may be realized to try to explain why the signal exhibits such non-Gaussianity (or some late time effect such as lensing). However, we shall soon find out that the PTA/CMB analogy ends here; because in the CMB, the signal is larger compared to instrumental noise, whereas in PTAs, the GWB signal is subdominant to the intrinsic pulsar red noises Agazie et al. (2023b). This compels us to venture into the signatures of a Gaussian GWB in two-point function statistics, utilizing both the power spectrum and the spatial correlation in the signal. Referring to our PTA simulations with only one hundred pulsars, this has shown some promise of not only resolving the HD signal in the mean statistic, but most importantly also the signatures of Gaussianity in the cumulants of the two-point function.
@@
This work proceeds as follows. In Section II, we first revisit the meaning of a Gaussian stochastic field and a Gaussian GWB (Sections II.1-II.2). We follow this up by elaborating on aspects of the one-point function that manifests said Gaussianity of a GWB signal in PTA simulations (Sections II.3-II.4). However, as we have teased out already, in PTAs, the one-point function statistics is not going to be able to resolve the Gaussianity information in the cumulants because of the noise in pulsars are dominant compared to the signal. We then move to the two-point function which turns out to be able to potentially display the signatures of a Gaussian GWB (Section III). Appendix A dwells into an intricacy between the central limit theorem and large scale modes. Our codes and notebooks that can be used to reproduce the main results of this work can be found in GitHub.
II Gravitational wave background statistics
@@
We briefly revisit the notion of a Gaussian field (Section II.1), a Gaussian GWB (Section II.2) and study its implications for the one-point function statistics in PTA (Sections II.3-II.4).
II.1 Gaussian statistics
@@
A Gaussian stochastic field, $\Psi(x)$, is completely specified by a two-point function, $\langle\Psi(x)\Psi(x^{\prime})\rangle$ Weinberg (2008). The odd moments of the field vanishes,
$\langle\Psi(x)\rangle=\langle\Psi(x)\Psi(x^{\prime})\Psi(x^{\prime\prime})\rangle=\cdots=0\,,$ (1)
@@
and the even moments factorize into a product-sum of two-point functions, $\langle\Psi(x)\Psi(x^{\prime})\rangle$; the four-point function becomes
$\begin{split}&\langle\Psi_{1}\Psi_{2}\Psi_{3}\Psi_{4}\rangle=\langle\Psi_{1}\Psi_{2}\rangle\langle\Psi_{3}\Psi_{4}\rangle+\langle\Psi_{1}\Psi_{3}\rangle\langle\Psi_{2}\Psi_{4}\rangle+\langle\Psi_{1}\Psi_{4}\rangle\langle\Psi_{2}\Psi_{3}\rangle\,,\end{split}$ (2)
@@
where we write down $\Psi_{i}=\Psi(x_{i})$ for brevity. The six-point, eight-point, and higher-point functions follow suit. For this work, we shall find the following two-point identities particularly useful Srednicki (1993); Gangui and Perivolaropoulos (1995),
$\langle\Psi_{1}^{3}\Psi_{2}^{3}\rangle=9\langle\Psi_{1}^{2}\rangle\langle\Psi_{2}^{2}\rangle\langle\Psi_{1}\Psi_{2}\rangle+6\langle\Psi_{1}\Psi_{2}\rangle^{3}\,,$ (3)
@@
$\langle\Psi_{1}^{4}\Psi_{2}^{4}\rangle=9\langle\Psi_{1}^{2}\rangle^{2}\langle\Psi_{2}^{2}\rangle^{2}+72\langle\Psi_{1}^{2}\rangle\langle\Psi_{2}^{2}\rangle\langle\Psi_{1}\Psi_{2}\rangle^{2}+24\langle\Psi_{1}\Psi_{2}\rangle^{4}\,.$ (4)
@@
In a Gaussian field, all information in the higher-point functions and cumulants can be unpacked into the two-point function. Thus, the Gaussianity of a field can be probed by measuring its moments. In practice, an observation $O$ can be taken as $f\left[\Psi(x)\right]$ where $f$ is some functional of the field $\Psi(x)$. If $\Psi(x)$ is a Gaussian field, and $f$ is a linear functional, then the observation $O$ can be expected to inherit the Gaussianity of the field; that is, the statistics of $O$ will be completely specified by a two-point function $\langle O(x)O(x^{\prime})\rangle$, and its higher moments will factorize as (1-4) with $O(x)$ in the place of the field $\Psi(x)$.
@@
This is the idea behind the statistics and isotropy tests of the CMB Akrami et al. (2020), where $\Psi(x)\equiv\delta(x)$ are scalar density fluctuations, and $O(x)\equiv\Delta T(x)$ is the temperature. In PTA science, $\Psi(x)\equiv h(x)$ are gravitational waves, and $O(x)\equiv r(x)$ are pulsar timing residuals. There are practical limits to this analogy between the CMB and the GWB in GW detectors such as PTAs. For the CMB, the presence of power at all angular scales, and subdominant instrumental noise, enable one-point statistics to be measured to sufficient accuracy to test the Gaussianity of the signal. For PTAs, neither is true; loud pulsar noises are generally present and the GWB power is concentrated in large scale modes. We consider the implications of this in the following sections.
II.2 The Gaussian GWB
We consider an isotropic and Gaussian GWB Hellings and Downs (1983); Phinney (2001),
$\langle h\left(f,\hat{k}\right)h^{*}\left(f^{\prime},\hat{k}^{\prime}\right)\rangle={\cal P}(f)\delta\left(f-f^{\prime}\right)\delta\left(\hat{k}-\hat{k}^{\prime}\right)\,,$ (5)
@@ II.2 The Gaussian GWB
where $h\left(f,\hat{k}\right)$ are GW amplitudes, $f$ is a GW frequency with a unit wavevector $\hat{k}$, and ${\cal P}(f)$ is the power spectrum. Polarization indices in the power spectrum are suppressed for conciseness. The assumption of Gaussianity implies that all the higher point information and the cumulants of the field in the GWB can be described by ${\cal P}(f)$. Observational departures from Gaussianity can be probed through the cumulants of the relevant observable Akrami et al. (2020).
@@
In weak fields relevant to PTA, we can expect that the statistics of the GWB are going to be inherited by the pulsar timing residuals. This allows an analytical route to express the cumulants of the pulsar timing residuals’ one-point function, and the corresponding cosmic variances, solely in terms of the angular power spectrum Srednicki (1993); Gangui and Perivolaropoulos (1995). For our discussion, we would sometimes refer to the cosmic variance Roebber et al. (2016); Roebber and Holder (2017); Allen (2023); Bernardo and Ng (2022, 2023a); Wu et al. (2024) as the ensemble variance, and the distribution of the values in different simulations/universes to be an ensemble distribution.
@@
Throughout this work, we quantify the statistics using the sample mean (${\cal E}_{1}$), variance (${\cal V}_{2}$), skewness (${\cal S}_{3}$), and kurtosis (${\cal K}_{4}$) of a set of points/pixels on the sphere, $s\equiv\{s_{1},s_{2},\cdots,s_{N}\}$, as ${\cal E}_{1}[s]=\sum_{i}s_{i}/N$, ${\cal V}_{2}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{2}\right]$, ${\cal S}_{3}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{3}\right]$, and ${\cal K}_{4}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{4}\right]$ [^1] We are loosely interchanging the terms for the skewness and the kurtosis with the third and fourth centralized moments. Our results hold for either statistical description.. The validity of the one-point statistics following Srednicki (1993); Gangui and Perivolaropoulos (1995) can be tested by numerically simulating the ensemble distribution of the sample statistics. Our analysis has shown that the analogous results of Srednicki (1993); Gangui and Perivolaropoulos (1995) for PTA holds; however, since the ensemble distribution of the one-point sample statistics turns out to be generally non-Gaussian distributed due to the spatial correlation, the first two moments, the mean and the cosmic variance, are no longer able to give a faithful picture of the true distribution (Appendix A). We shall show that this is the case for both the GWB signal and red noise in PTA simulations, but with differing natures of induced non-Gaussianities.
II.3 Mock PTA data
@@
We generate our PTA simulations following Babak et al. (2024). This gives timing data for each pulsar in the form
$r(t,\hat{e}_{a})=\sum_{k}\left(\alpha_{ak}\sin(\omega_{k}t)+\beta_{ak}\cos(\omega_{k}t)\right)\,,$ (6)
@@
For red noise, the frequency components are drawn from a Gaussian distribution, $\alpha_{ak},\beta_{bk}\sim{\cal N}\left(0,S_{ab}(f_{k})\Delta f_{k}\right)$ where $S_{ab}(f)=\delta_{ab}{\cal S}^{\rm N}(f)$ is the noise power spectrum and $\Delta f_{k}=f_{k+1}-f_{k}$. We consider a simple power law for the red noise spectrum. In this case, the noise components are uncorrelated across different pulsars. For the GWB, the data is obtained by $\alpha_{ak},\beta_{bk}\sim{\cal N}\left(0,S_{ab}(f_{k})\Delta f_{k}\right)$ where $S_{ab}(f)=\Gamma_{ab}{\cal S}^{\rm C}(f)$, $\Gamma_{ab}$ is the HD correlation, and ${\cal S}^{\rm C}(f)$ is the power spectrum of circular SMBHBs. In this case, the GWB signal is viewed as a common spectrum that is spatially correlated across pulsars, in accordance with the HD curve.
II.4 One-point statistics
For the GWB, we consider an amplitude $A_{\rm gw}=2.4\times 10^{-15}$ and spectral index $\gamma_{\rm gw}=13/3$ corresponding to circular SMBHBs Phinney (2001), along with the HD correlation. For the red noise (RN), considered as independent Gaussian random processes with a power law spectrum, we uniformly draw the amplitudes, $\log_{10}A_{\rm rn}={\cal U}(-17,-13)$, and spectral indices, $\gamma_{\rm rn}={\cal U}(2,6)$, for each pulsar.

### v2

@@ 1 Introduction
The first direct detection of a gravitational wave (GW) has ushered in a new era in astronomy [1]. Hundreds of GW events have now been observed from Solar mass compact objects [2], each one carrying important information about the strong gravity regime [3]. The next breakthrough in GW astronomy is the detection of a gravitational wave background (GWB) [4, 5, 6, 7, 8, 9, 10], coming from a superposition of GWs from sources that are too weak to be resolved individually. Efforts are ongoing in order to resolve the GWB in the sub-kiloherz band using ground-based GW detectors [11, 12, 13] and in the millihertz band in future space-based GW observations [14, 15, 16, 17]. Recently, pulsar timing array (PTA) collaborations have reported a compelling evidence in favor of the presence of a nanohertz stochastic GWB [18, 19, 20, 21, 22], consistent with a common-spectrum process across pulsars with the signature Hellings and Downs (HD) correlation [23]. This signal is traditionally interpreted as coming from a population of supermassive black hole binaries (SMBHB) [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. However, the PTAs’ observation also allows for an interesting set of sources of the nanohertz GWB, rooted cosmologically [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. A challenge to the field now is to identify the signatures in the observation that can distinguish the different sources of a GWB, ultimately settling the issue of whether an observed GWB signal is astrophysical or cosmological.
@@
Different sources of a GWB produce the same HD spatial correlation, only with a different spectrum. A logical step to determine the source of the the signal is thus by calculating the expected spectrum for each source and computing the respective evidences relative to the data. This approach has been widely adopted by the PTA community, but has not ruled out any significant possibilities, given the present data [52, 53, 54]. The field is optimistic that forthcoming iterations of the data [55, 56, 57, 58], with more and longer timed pulsars, may eventually enable PTAs to contribute to the source debate. This poses the question whether there are further untapped aspects of the observation that may be able to help out in distinguishing the source of an observed GWB signal, not only in PTA but in general. This brings us to the theme of this work which is testing the Gaussianity of a GWB.
@@
In this work, we bare down signatures of the GWB signal in the one- and two-point functions’ cumulants, referring to PTA simulations [59] as a proof-of-concept. In particular, a Gaussian GWB requires that all the higher-point information and the cumulants of the observable reduce to products of the power spectrum, and the spatial correlation. The relation this has on the source is that the level of Gaussianity that is expected to be carried by the signal is different for astrophysical/SMBHBs and cosmological sources, with the latter expected to display a higher degree of Gaussianity. Testing the Gaussianity of the signal thus boils down to measuring the cumulants of a relevant observable, and checking if there is additional information in the cumulants that is not already present in the power spectrum.
@@
This direction we embark on is reminiscent of tests of isotropy and statistics in the cosmic microwave background (CMB) [60]. In CMB analysis, the one-point information is constrained to assess if the data shows any significant departures from Gaussianity. If there is, then a primordial theory may be realized to try to explain why the signal exhibits such non-Gaussianity (or some late time effect such as lensing). However, we shall soon find out that the PTA/CMB analogy ends here; because in the CMB, the signal is larger compared to instrumental noise, whereas in PTAs, the GWB signal is subdominant to the intrinsic pulsar red noises [61, 62, 63]. This compels us to venture into the signatures of a Gaussian GWB in two-point function statistics, utilizing both the power spectrum and the spatial correlation in the signal. Referring to our PTA simulations with only one hundred pulsars, this has shown some promise of not only resolving the HD signal in the mean statistic, but most importantly also the signatures of Gaussianity in the cumulants of the two-point function.
@@
This work proceeds as follows. In Section 2, we first revisit the meaning of a Gaussian stochastic field and a Gaussian GWB (Sections 2.1-2.2). We follow this up by elaborating on aspects of the one-point function that manifests said Gaussianity of a GWB signal in PTA simulations (Sections 2.3-2.4). However, as we have teased out already, in PTAs, the one-point function statistics is not going to be able to resolve the Gaussianity information in the cumulants because of the noise in pulsars are dominant compared to the signal. We then move to the two-point function which turns out to be able to potentially display the signatures of a Gaussian GWB (Section 3). Appendix A dwells into an intricacy between the central limit theorem and large scale modes. Our codes and notebooks that can be used to reproduce the main results of this work can be found in GitHub.
2 Gravitational wave background statistics
@@
We briefly revisit the notion of a Gaussian field (Section 2.1), a Gaussian GWB (Section 2.2) and study its implications for the one-point function statistics in PTA (Sections 2.3-2.4).
2.1 Gaussian statistics
@@
A Gaussian stochastic field, $\Psi(x)$, is completely specified by a two-point function, $\langle\Psi(x)\Psi(x^{\prime})\rangle$, a one-point (that can always be shifted to zero), and higher-point functions that admit the relations described below [64]. The odd moments of the field vanishes,
$\langle\Psi(x)\rangle=\langle\Psi(x)\Psi(x^{\prime})\Psi(x^{\prime\prime})\rangle=\cdots=0\,,$ (2.1)
@@
and the even moments factorize into a product-sum of two-point functions, $\langle\Psi(x)\Psi(x^{\prime})\rangle$; the four-point function becomes [65]
$\begin{split}&\langle\Psi_{1}\Psi_{2}\Psi_{3}\Psi_{4}\rangle=\langle\Psi_{1}\Psi_{2}\rangle\langle\Psi_{3}\Psi_{4}\rangle+\langle\Psi_{1}\Psi_{3}\rangle\langle\Psi_{2}\Psi_{4}\rangle+\langle\Psi_{1}\Psi_{4}\rangle\langle\Psi_{2}\Psi_{3}\rangle\,,\end{split}$ (2.2)
@@
where we write down $\Psi_{i}=\Psi(x_{i})$ for brevity. The six-point, eight-point, and higher-point functions follow suit. For this work, we shall find the following two-point identities particularly useful [66, 67],
$\langle\Psi_{1}^{3}\Psi_{2}^{3}\rangle=9\langle\Psi_{1}^{2}\rangle\langle\Psi_{2}^{2}\rangle\langle\Psi_{1}\Psi_{2}\rangle+6\langle\Psi_{1}\Psi_{2}\rangle^{3}\,,$ (2.3)
@@
$\langle\Psi_{1}^{4}\Psi_{2}^{4}\rangle=9\langle\Psi_{1}^{2}\rangle^{2}\langle\Psi_{2}^{2}\rangle^{2}+72\langle\Psi_{1}^{2}\rangle\langle\Psi_{2}^{2}\rangle\langle\Psi_{1}\Psi_{2}\rangle^{2}+24\langle\Psi_{1}\Psi_{2}\rangle^{4}\,.$ (2.4)
@@
In a Gaussian field, all information in the higher-point functions and cumulants can be unpacked into the two-point function. Thus, the Gaussianity of a field can be probed by measuring its moments. In practice, an observation $O$ can be taken as $f\left[\Psi(x)\right]$ where $f$ is some functional of the field $\Psi(x)$. If $\Psi(x)$ is a Gaussian field, and $f$ is a linear functional, then the observation $O$ can be expected to inherit the Gaussianity of the field; that is, the statistics of $O$ will be completely specified by a two-point function $\langle O(x)O(x^{\prime})\rangle$, and its higher moments will factorize as (2.1-2.4) with $O(x)$ in the place of the field $\Psi(x)$.
@@
This is the idea behind the statistics and isotropy tests of the CMB [60], where $\Psi(x)\equiv\delta(x)$ are scalar density fluctuations, and $O(x)\equiv\Delta T(x)$ is the temperature. In PTA science, $\Psi(x)\equiv h(x)$ are gravitational waves, and $O(x)\equiv r(x)$ are pulsar timing residuals. There are practical limits to this analogy between the CMB and the GWB in GW detectors such as PTAs. For the CMB, the presence of power at all angular scales, and subdominant instrumental noise, enable one-point statistics to be measured to sufficient accuracy to test the Gaussianity of the signal. For PTAs, neither is true; loud pulsar noises are generally present and the GWB power is concentrated in large scale modes. We consider the implications of this in the following sections.
2.2 The Gaussian GWB
We consider a stationary, isotropic and Gaussian GWB [23, 26],
$\langle h_{{\color[rgb]{0,0,1}A}}\left(f,\hat{k}\right)h_{{\color[rgb]{0,0,1}A^{\prime}}}^{*}\left(f^{\prime},\hat{k}^{\prime}\right)\rangle={\cal P}(f){{\color[rgb]{0,0,1}\delta_{AA^{\prime}}}}\delta\left(f-f^{\prime}\right)\delta\left(\hat{k}-\hat{k}^{\prime}\right)\,,$ (2.5)
@@ 2.2 The Gaussian GWB
where $h_{{\color[rgb]{0,0,1}A}}\left(f,\hat{k}\right)$ are GW amplitudes, $A=+,\times$ are GW polarization indices, $f$ is a GW frequency with a unit wavevector $\hat{k}$, and ${\cal P}(f)$ is the power spectrum. The assumption of Gaussianity implies that all the higher point information and the cumulants of the field in the GWB can be described by ${\cal P}(f)$. Observational departures from Gaussianity can be probed through the cumulants of the relevant observable [60].
@@
In weak fields relevant to PTA, we can expect that the statistics of the GWB are going to be inherited by the pulsar timing residuals. This allows an analytical route to express the cumulants of the pulsar timing residuals’ one-point function, and the corresponding cosmic variances, solely in terms of the angular power spectrum [66, 67]. For our discussion, we would sometimes refer to the cosmic variance [68, 69, 70, 71, 72, 73] as the ensemble variance, and the distribution of the values in different simulations/universes to be an ensemble distribution.
@@
Throughout this work, we quantify the statistics using the sample mean (${\cal E}_{1}$), variance (${\cal V}_{2}$), skewness (${\cal S}_{3}$), and kurtosis (${\cal K}_{4}$) of a set of points/pixels on the sphere, $s\equiv\{s_{1},s_{2},\cdots,s_{N}\}$, as ${\cal E}_{1}[s]=\sum_{i}s_{i}/N$, ${\cal V}_{2}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{2}\right]$, ${\cal S}_{3}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{3}\right]$, and ${\cal K}_{4}[s]={\cal E}_{1}\left[\left(s-{\cal E}_{1}[s]\right)^{4}\right]$ [^2] We are loosely interchanging the terms for the skewness and the kurtosis with the third and fourth centralized moments. Our results hold for either statistical description.. The quantity $s$ represents the pulsar timing residuals (sum $i$ over $n_{\rm psrs}$ pulsars, Section 2.4) or the residual-pair product (sum $i$ over $n_{\rm psrs}(n_{\rm psrs}-1)/2$ pulsar pairs, Section 3). The validity of the one-point statistics following [66, 67] can be tested by numerically simulating the ensemble distribution of the sample statistics. Our analysis has shown that the analogous results of [66, 67] for PTA holds; however, since the ensemble distribution of the one-point sample statistics turns out to be generally non-Gaussian distributed due to the spatial correlation, the first two moments, the mean and the cosmic variance, are no longer able to give a faithful picture of the true distribution (Appendix A). We shall show that this is the case for both the GWB signal and red noise in PTA simulations, but with differing natures of induced non-Gaussianities.
2.3 Mock PTA data
@@
We generate our PTA simulations following [59]. This gives timing data for each pulsar in the form
$r(t,\hat{e}_{a})=\sum_{k}\left(\alpha_{ak}\sin(\omega_{k}t)+\beta_{ak}\cos(\omega_{k}t)\right)\,,$ (2.6)
@@
For red noise, the frequency components are drawn from a Gaussian distribution, $\alpha_{ak},\beta_{bk}\sim{\cal N}\left(0,S_{ab}(f_{k})\Delta f_{k}\right)$ where $S_{ab}(f)=\delta_{ab}{\cal S}^{\rm N}(f)$ is the noise power spectrum and $\Delta f_{k}=f_{k+1}-f_{k}$ [74]. Note that we use subscripts $a,b$ to label pulsars and $k$ for frequency bins. We consider a simple power law for the red noise spectrum. In this case, the noise components are uncorrelated across different pulsars. For the GWB, the data is obtained by $\alpha_{ak},\beta_{bk}\sim{\cal N}\left(0,S_{ab}(f_{k})\Delta f_{k}\right)$ where $S_{ab}(f)=\Gamma_{ab}{\cal S}^{\rm C}(f)$, $\Gamma_{ab}$ is the HD correlation, and ${\cal S}^{\rm C}(f)$ is the power spectrum of circular SMBHBs. In this case, the GWB signal is viewed as a common spectrum that is spatially correlated across pulsars, in accordance with the HD curve.
2.4 One-point statistics
For the GWB, we consider an amplitude $A_{\rm gw}=2.4\times 10^{-15}$ and spectral index $\gamma_{\rm gw}=13/3$ corresponding to circular SMBHBs [26], along with the HD correlation. For the red noise (RN), considered as independent Gaussian random processes with a power law spectrum, we uniformly draw the amplitudes, $\log_{10}A_{\rm rn}={\cal U}(-17,-13)$, and spectral indices, $\gamma_{\rm rn}={\cal U}(2,6)$, for each pulsar. This gives a time series data $r(t,\hat{e}_{a})$, for $a=1,2,\cdots n_{\rm psrs}$ pulsars, and the Fourier bins $\alpha_{ak},\beta_{ak}$ (via a standard discrete Fourier transform routine, see Eq. (2.6)). Then, the sample statistics of $\alpha_{ak},\beta_{ak}$ are computed over the simulated PTA (one realisation); average over pulsars in the PTA. The entire process is repeated numerous times with the same signal and noise priors, each realisation turning in values of the sample statistics representative of an underlying ensemble PDF/cosmic variance, as shown in Figure 1.